# Effect of Off-diagonal NSI Parameters on Entanglement Measurements in Neutrino Oscillations

### v1

@@ V.2 Analysis based on quantum entanglement measures
Figure 6 illustrates how the NSI parameter, $\epsilon_{e\mu}$ affects EOF, Concurrence, and Negativity. In each panel, the red curve corresponds to the standard oscillation scenario with no NSI present. The green curve shows the effect of a non-zero real part, specifically $\epsilon_{e\mu}=0.15$ with the associated phase $\phi_{e\mu}=0$. The blue curve includes both the real and imaginary components of $\epsilon_{e\mu}$, i.e., the full complex parameter. The shaded region represents the $\nu_{\mu}$ flux in arbitrary units. The left, middle, and right panels of Fig. 6 respectively depict the variation of EOF, Concurrence, and Negativity due to the presence of $\epsilon_{e\mu}$. For these plots, the standard oscillation parameters are taken from Table 1, while the NSI values are from Table 2. Upper row of Fig. 6 shows the result when $\delta_{CP}=212^{\circ}$ whereas lower row represents the result with $\delta_{CP}=0^{\circ}$.
@@
To understand the influence of the $\epsilon_{e\mu}$ parameter on the Entanglement of Formation (EOF), we refer to the leftmost panel of the upper row in Fig. 6. Around the DUNE flux oscillation peak, approximately at 2.5 GeV, a clear deviation of the blue curve from the standard oscillation curve (red) is observed. When both the real and imaginary components of $\epsilon_{e\mu}$ are included, the EOF value decreases at this oscillation peak. In contrast, if only the real part of $\epsilon_{e\mu}$ is considered, there is a slight enhancement in the EOF amplitude at the same energy compared to the standard case. This behavior continues consistently up to around 5 GeV, beyond which all three curves converge and exhibit negligible differences. In the lower row of the leftmost panel, with $\delta_{CP}=0^{\circ}$, the EOF displays a different trend. Here, the green curve diverges significantly from both the red and blue curves, particularly beyond 4 GeV, where the difference between the green and red curves becomes more pronounced.
@@
To explain the nature of both the panels, we need to see the numerical expression. The expression of EOF in terms of oscillation probabilities are given in Eq. III.1. In this study, we mainly focus on the accelerator-based long-baseline experiment DUNE. For the DUNE experimental setup, the experiment is sensitive to measure the appearance and disappearance probabilities of muon electron neutrinos. Thus, we need to express Eq. III.1 in terms of $P_{\mu e}$ and $P_{\mu\mu}$. The modified form of the equation III.1 (considering $\nu_{\mu}$ as initial neutrino flavor) in terms of $P_{\mu e}$ and $P_{\mu\mu}$ is given by:
@@
×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times\log_{2}(1-P_{\mu e})+(1-P_{\mu\mu})\log_{2}(1-P_{\mu\mu})+(P_{\mu\mu}+P_{\mu e})\log_{2}(P_{\mu\mu}+P_{\mu e})\Big{]}.$ (24)
From Eq. V.2, it is evident that the EOF depends equally on both the appearance probability $P_{\mu e}$ and the disappearance probability $P_{\mu\mu}$. However, as shown in Eq. V.1 and Fig. 4, the influence of the NSI parameter $\epsilon_{e\mu}$ on the disappearance channel is minimal. Therefore, the behavior of EOF is primarily governed by the appearance probability in presence of the NSI parameter $\epsilon_{e\mu}$, as described in Eq. V.1 and illustrated in Fig. 1.
@@
Next, we turn our attention to understanding how $\epsilon_{e\mu}$ influences the other two entanglement measures: Concurrence and Negativity.
The simplified form of Concurrence (Eq. 12) in terms of electron neutrino appearance and muon neutrino disappearance probabilities as,
@@
Figure 7 illustrates the impact of the NSI parameter $\epsilon_{e\tau}$ on the three entanglement measures: EOF, Concurrence, and Negativity. The color scheme remains consistent with the previous figure; red denotes the standard oscillation scenario, green corresponds to a non-zero real value of $\epsilon_{e\tau}$ with phase $\phi_{e\tau}=0$ rad, and blue represents the case where both real and imaginary parts of $\epsilon_{e\tau}$, along with its associated phase $\phi_{e\tau}$, are considered. The shaded area corresponds to the DUNE $\nu_{\mu}$ neutrino flux in arbitrary units.
In the upper row with $\delta_{CP}=212^{\circ}$, all three measures show that the green curve lies below both the red and blue curves at the oscillation peak. At higher neutrino energies, the three curves converge for EOF and Concurrence, indicating negligible differences. However, in the case of Negativity, the curves remain distinct even at higher energies. For $\delta_{CP}=0^{\circ}$, shown in the lower row, the blue curve lies below the red and green curves at the oscillation maximum across all three entanglement measures. As the energy increases, the curves tend to overlap for EOF and Concurrence, while they remain clearly separated for Negativity. The overall behavior observed in Fig. 7 can be attributed solely to the effect of $\epsilon_{e\tau}$ on the appearance probability, as the disappearance probability $P_{\ mu\mu}$ is unaffected by this parameter.
@@
In summary, for long-baseline accelerator neutrino experiments, the influence of off-diagonal NSI parameters on entanglement measures is primarily governed by their impact on the appearance and disappearance channels. Specifically, the parameters $\epsilon_{e\mu}$ and $\epsilon_{e\tau}$ predominantly affect the appearance probability $P_{\mu e}$, while $\epsilon_{\mu\tau}$ modifies the disappearance probability $P_{\mu\mu}$.
@@
Additionally, similar patterns in the entanglement measures; EOF, Concurrence, and Negativity are observed in future long-baseline experiments such as P2SO when these off-diagonal NSI parameters are introduced.
VI Conclusion
@@
In this study, we investigated the impact of off-diagonal non-standard interaction (NSI) parameters on quantum entanglement within the three-flavor neutrino oscillation framework. By reformulating three commonly used entanglement measures; Entanglement of Formation (EOF), Concurrence, and Negativity in terms of oscillation probabilities, we analyzed how quantum correlations are modified by the presence of the off-diagonal NSI parameters $\epsilon_{e\mu}$, $\epsilon_{e\tau}$, and $\epsilon_{\mu\tau}$, including the influence of their complex phases.
@@
Focusing on the DUNE experimental setup, we studied how both the magnitude and phase of these parameters affect the energy-dependent behavior of the entanglement measures. Our findings indicate that NSI effects are most prominent at lower neutrino energies, specifically at the oscillation maximum position. Among the three measures, Negativity consistently shows the most pronounced sensitivity to NSI, clearly distinguishing scenarios with non-zero phases from standard oscillations, even at higher energies. Moreover, we found that the entanglement behavior is predominantly governed by the oscillation (or survival) probabilities: the appearance channel plays the dominant role when $\epsilon_{e\mu}$ or $\epsilon_{e\tau}$ is considered individually, while the disappearance channel becomes central in the case of $\epsilon_{\mu\tau}$.

### v2

@@ V.3 Analysis based on quantum entanglement measures
Figure 7 illustrates how the NSI parameter, $\epsilon_{e\mu}$ affects EOF, Concurrence, and Negativity. In each panel, the red curve corresponds to the standard oscillation scenario with no NSI present. The green curve shows the effect of a non-zero real part, specifically $\left|\epsilon_{e\mu}\right|$ = 0.15 with the associated phase $\phi_{e\mu}=0$. The blue curve includes both the real and imaginary components of $\epsilon_{e\mu}$, i.e., the full complex parameter. The shaded region represents the $\nu_{\mu}$ flux in arbitrary units. The left, middle, and right panels of Fig. 7 respectively depict the variation of EOF, Concurrence, and Negativity due to the presence of $\epsilon_{e\mu}$. For these plots, the standard oscillation parameters are taken from Table 1, while the NSI values are from Table 2. Upper row of Fig. 7 shows the result when $\delta_{CP}=212^{\circ}$ whereas lower row represents the result with $\delta_{CP}=0^{\circ}$.
@@
To understand the influence of the $\epsilon_{e\mu}$ parameter on the Entanglement of Formation (EOF), we refer to the leftmost panel of the upper row in Fig. 7. A clear deviation of the blue curve from the standard oscillation curve (red) is observed near the DUNE flux oscillation peak approximately at 2.5 GeV, corresponding to an approximately 28% deviation from the SO scenario for $\delta_{CP}=212^{\circ}$. When both the real and imaginary components of $\epsilon_{e\mu}$ are included, the EOF value decreases at this oscillation peak. In contrast, if only the real part of $\epsilon_{e\mu}$ is considered, there is a slight enhancement in the EOF amplitude at the same energy compared to the standard case. This behavior continues consistently up to around 5 GeV, beyond which all three curves converge and exhibit negligible differences. In the lower row of the leftmost panel, with $\delta_{CP}=0^{\circ}$, the EOF displays a different trend. Here, the green curve diverges significantly from both the red and blue curves, particularly beyond 4 GeV, where the difference between the green and red curves becomes more pronounced, reaching approximately 18%. A similar qualitative behavior is observed for Concurrence. For $\delta_{CP}=212^{\circ}$, the red curve deviates from the blue curve by approximately 38% at $E\approx$ 2.5 GeV. In contrast, for $\delta_{CP}=0^{\circ}$, the green curve deviates from the red curve by about 6% at energies above 4 GeV. For Negativity, in the upper-left panel, the blue curve can be clearly distinguished from the red curve at $E$ approx 2.5 GeV, with a deviation of approximately 16%. However, for $\delta_{CP}=0^{\circ}$, at energies above 4 GeV, all three curves (red, blue, and green) become clearly distinguishable, with the blue curve deviating from the red by about 9% and the green curve by approximately 31%.
@@
To explain the nature of both the panels, we need to see the numerical expression. The expression of EOF in terms of oscillation probabilities is given in Eq. III.1. In this study, we mainly focus on the accelerator-based long-baseline experiment DUNE. For the DUNE experimental setup, the experiment is sensitive to measure the appearance and disappearance probabilities of muon electron neutrinos. Thus, we need to express Eq. III.1 in terms of $P_{\mu e}$ and $P_{\mu\mu}$. The modified form of the Eq. III.1 (considering $\nu_{\mu}$ as initial neutrino flavor) in terms of $P_{\mu e}$ and $P_{\mu\mu}$ is given by:
@@
$~~~~~~~~~~\times\log_{2}(1-P_{\mu e})+(1-P_{\mu\mu})\log_{2}(1-P_{\mu\mu})+(P_{\mu\mu}+P_{\mu e})\log_{2}(P_{\mu\mu}+P_{\mu e})\Big].$ (31)
From Eq. V.3, it is evident that the EOF depends equally on both the appearance probability $P_{\mu e}$ and the disappearance probability $P_{\mu\mu}$. However, as shown in Eq. V.1 and Fig. 4, the influence of the NSI parameter $\epsilon_{e\mu}$ on the disappearance channel is minimal. Therefore, the behavior of EOF is primarily governed by the appearance probability in the presence of the NSI parameter $\epsilon_{e\mu}$, as described in Eq. V.1 and illustrated in Fig. 1.
@@
Next, we turn our attention to understanding how $\epsilon_{e\mu}$ influences the other two entanglement measures: Concurrence and Negativity. The simplified form of Concurrence (Eq. 19) in terms of electron neutrino appearance and muon neutrino disappearance probabilities as,
@@
Figure 8 illustrates the impact of the NSI parameter $\epsilon_{e\tau}$ on the three entanglement measures: EOF, Concurrence, and Negativity. The color scheme remains consistent with the previous figure; red denotes the standard oscillation scenario, green corresponds to a non-zero real value of $\epsilon_{e\tau}$ with phase $\phi_{e\tau}=0$, and blue represents the case where both real and imaginary parts of $\epsilon_{e\tau}$, along with its associated phase $\phi_{e\tau}$, are considered. The shaded area corresponds to the DUNE $\nu_{\mu}$ neutrino flux in arbitrary units.
In the upper row with $\delta_{CP}=212^{\circ}$, all three measures show that the green curve lies below both the red and blue curves at the oscillation peak, with suppressions of approximately 50% for EOF, $\approx$ 62% for Concurrence, and approx 33% for Negativity. At higher neutrino energies, the three curves converge for EOF and Concurrence, indicating negligible differences. However, in the case of Negativity, the curves remain distinct even at higher energies. For energies above 4 GeV, the blue curve shows a modest enhancement of about 2% relative to the red curve, whereas the green curve exhibits a significant suppression of approximately 14% compared to the red curve. For $\delta_{CP}=0^{\circ}$, shown in the lower row, the blue curve lies below the red and green curves at the oscillation maximum across all three entanglement measures. At this energy, relative to the red curve, the green curve exhibits an enhancement of about 36% in EOF, $\approx$ 60% in concurrence, and $\approx$ 30% in negativity. In contrast, the blue curve shows a suppression of $\approx$ 22% for EOF and approx 33% for concurrence, while for negativity only a mild suppression is observed. As the energy increases, the curves tend to overlap for EOF and Concurrence, while they remain clearly separated for Negativity. The overall behavior observed in Fig. 8 can be attributed solely to the effect of $\epsilon_{e\tau}$ on the appearance probability, as the disappearance probability $P_{\mu\mu}$ is unaffected by this parameter.
Figure 10: The entanglement measures: EOF (top), Concurrence (middle), and Negativity (bottom) are shown in the plane of the neutrino energy E (in GeV) and the CP-violating phase $\delta_{CP}$ (in degree) for the SO (left), $\left|\epsilon_{e\mu}\right|,\phi_{e\mu}=0$ (middle)
, and $\left|\epsilon_{e\mu}\right|,\phi_{e\mu}$ (right) scenarios for DUNE experimental set up.
@@
In summary, for long-baseline accelerator neutrino experiments, the influence of off-diagonal NSI parameters on entanglement measures is primarily governed by their impact on the appearance and disappearance channels. Specifically, the NSI parameters $\epsilon_{e\mu}$ and $\epsilon_{e\tau}$ predominantly affect the appearance probability $P_{\mu e}$, while $\epsilon_{\mu\tau}$ modifies the disappearance probability $P_{\mu\mu}$.
@@
Additionally, we have checked that similar patterns in the entanglement measures; EOF, Concurrence, and Negativity are observed in future long-baseline experiments such as P2SO when these off-diagonal NSI parameters are introduced.
Figure 11: For the DUNE experiment, the entanglement measures: EOF (top), Concurrence (middle), and Negativity (bottom) are plotted in the (E - $\delta_{CP}$) plane for the SO (Left), $\epsilon_{e\tau},\phi_{e\tau}=0$ (middle) and, $\epsilon_{e\tau},\phi_{e\tau}$ (right) scenarios.
V.4 Dependency of $\delta_{CP}$ and energy on entanglement measures
In this subsection, we mainly focus on the dependency of $\delta_{CP}$ and neutrino energy on EOF, Concurrence and Negativity in the absence and presence of NSI parameters. In Fig. 10, we show the impact of off-diagonal NSI parameter $\left|\epsilon_{e\mu}\right|$ and its associated complex phase $\phi_{e\mu}$ on the entanglement measures as a function of neutrino energy and the CP-violating phase $\delta_{CP}$ for the DUNE experimental set-up. Each row in the figure corresponds to one entanglement measure, while the three columns represent different scenarios: the standard oscillation (SO), NSI with a real valued $\epsilon_{e\mu}$ $(\phi_{e\mu}=0)$, and NSI with both magnitude and complex phase of $\epsilon_{e\mu}$ $(\phi_{e\mu}\neq 0)$. It is observed that in the energy range of [1.52][1.5-2] GeV, for all scenarios, EOF achieved the maximum value ($\sim 1.2$) for all the values of $\delta_{CP}$. Further in the energy range of around [46][4-6] GeV, for SO and $\left|\epsilon_{e\mu}\right|$ with non-zero $\phi_{e\mu}$ scenarios, the EOF shows maximum and for the scenario with $\phi_{e\mu}=0$, the energy range for maximum EOF value is around [47][4-7] GeV for all values of $\delta_{CP}$ phase. At 2.5 GeV, which corresponds to the oscillation flux peak in DUNE, the EOF is observed to be relatively low. Similarly, Concurrence nearly mirrors the behavior of EOF, attaining a maximum value of about 1.2, at the energy range of [0.52][0.5-2] GeV and [46][4-6] GeV for SO and the condition when the magnitude and complex phase of NSI, both are present. The energy range for maximum Concurrence with the case of non-zero $\left|\epsilon_{e\mu}\right|$ and zero $\phi_{e\mu}$ is little broader, around [46.5][4-6.5] GeV. At the oscillation peak of the DUNE setup, similar to EOF, the Concurrence value shows nearly zero for almost all the energy range of $\delta_{CP}$.
Unlike EOF and Concurrence, Negativity shows quite different behavior. Lower row of Fig. 10, shows the variation of negativity with neutrino energy and CP-violating phase for SO (leftmost panel), presence of non-zero $\left|\epsilon_{e\mu}\right|$ and zero $\phi_{e\mu}$ (middle panel) and non-zero values of $\left|\epsilon_{e\mu}\right|$ and $\phi_{e\mu}$ (rightmost panel). For each panel, the maximum negativity can be seen only at the energy range of [00.5][0-0.5] GeV with the full range of $\delta_{CP}$. For the SO scenario, the value of negativity reaches a maximum in the energy window of [1.52][1.5-2] GeV, and this occurs over a distinct range of $50^{\circ}\leq\delta_{CP}\leq 225^{\circ}$ and in the case of the NSI $\epsilon_{e\mu}$ scenario without and with an associated phase $\phi_{e\mu}$, the maximum value of negativity observed within $50^{\circ}\leq\delta_{CP}\leq 200^{\circ}$ and $100^{\circ}\leq\delta_{CP}\leq 200^{\circ}$ respectively. In the energy range of [46][4-6] GeV, the negativity shows clear differentiation between $\delta_{CP}=0^{\circ}$ and $212^{\circ}$ in the SO scenario. For the NSI scenario with an associated complex phase $\phi_{e\mu}$, this distinctive behavior persists and even extends over a broader energy range of about [48][4-8] GeV, indicating an enhanced phase dependent response in the presence of NSI effects. At 2.5 GeV, negativity shows nearly zero value for all the three conditions.
In Fig. 11, three entanglement measures are shown in the plane of the neutrino energy $E$ and the CP-violating phase $\delta_{CP}$ for three scenarios: the SO, the presence of an off-diagonal NSI parameter $\left|\epsilon_{e\tau}\right|$ with a vanishing complex phase $(\phi_{e\tau}=0)$, and a scenario where both the magnitude and complex phase of $\epsilon_{e\tau}$ are non-zero $(\phi_{e\tau}\neq 0)$. The figure is organized such that each row displays one of the entanglement measures, while the columns compare three scenarios: the first shows standard oscillation, the second includes NSI with a real-valued $\epsilon_{e\tau}$, and the third incorporates both the magnitude and complex phase of $\epsilon_{e\tau}$. In the SO case, the EOF reaches its maximum value of approx 1.2 within the energy ranges of [1.52][1.5-2] GeV and [46][4-6] GeV, consistently across all values of $\delta_{CP}$. In contrast, for the NSI scenario with a vanishing complex phase $(\phi_{e\tau}=0)$, the maximum EOF occurs within the energy range [1.52][1.5-2] GeV and for $\delta_{CP}$ values around $50^{\circ}-250^{\circ}]$. Similarly, in the case of the NSI with non-vanishing complex phase scenario, the maximum EOF is also observed in the same energy range, but for $\delta_{CP}$ values approximately between 25 and 225. However, at this energy range, the EOF shows different values for $\delta_{CP}$, which is equal to 0 and 212, indicating a dependence on the CP-violating phase both in the presence and absence of the NSI phase. For the Concurrence, all the scenarios attain their maximum value, 1.2 in the energy ranges of [1.52][1.5-2] GeV and [46][4-6] GeV, consistently across the entire range of CP-violating phases, $\delta_{CP}$. However, negativity reaches its maximum value of 0.35 within the energy range of [1.52][1.5-2] GeV for all scenarios, though the corresponding ranges of the CP-violating phase $\delta_{CP}$ differ. Specifically, the maximum occurs for $\delta_{CP}$ values in the ranges of $120^{\circ}-160^{\circ}]$ for the SO case, approx $40^{\circ}-260^{\circ}]$ for the NSI scenario with vanishing complex phase, and $\approx[0^{\circ}-240^{\circ}]$ for the NSI scenario with non-vanishing complex phase.
Figure 12: EOF (top), Concurrence (middle), and Negativity (bottom) are shown in the plane of the neutrino energy E (in GeV) and the CP-violating phase $\delta_{CP}$ (in degree) for the SO (Left), $\epsilon_{\mu\tau},\phi_{\mu\tau}=0$ (middle), and $\epsilon_{\mu\tau},\phi_{\mu\tau}$ (right) scenarios for DUNE.
Figure 12 depicts the impact of NSI involving the $\epsilon_{\mu\tau}$ and their associated phase on entanglement measures, analyzed using three distinct measures plotted as functions of neutrino energy $E$ and the CP-violating phase $\delta_{CP}$. The analysis considers three separate scenarios: the SO without NSI, the presence of a non-zero $\epsilon_{\mu\tau}$ with zero complex phase, and a situation with both the magnitude and phase of $\epsilon_{\mu\tau}$. The EOF exhibits a pattern for the SO scenario reaching its maximum within the energy intervals of [1.52][1.5-2] GeV and [46][4-6] GeV across all values of the CP-violating phase $\delta_{CP}$, whereas for the NSI case with a vanishing complex phase the maximum shifted into higher energy range, around [5.59][5.5-9] GeV. However, in the NSI scenario with a non-vanishing complex phase, the peak of EOF is attained in the energy range [45][4-5] GeV for the $\delta_{CP}$ values from 0 to 180 and 300 to 360. Notably, for this case, a distinction is observed between the EOF values at $\delta_{CP}$= 0 and 212, highlighting the phase sensitivity even in the absence of a complex NSI phase. A comparable trend is observed for Concurrence across all three scenarios, SO, NSI with vanishing phase, and NSI with non-vanishing phase, with the maximum occurring in the specific energy ranges, largely independent of $\delta_{CP}$. However, a distinct minimum region appears for Concurrence in the $\delta_{CP}$ range of 190 to 350 for energy nearly 3 GeV, with noticeable differences between the values at $\delta_{CP}$= 0 and 212 for the NSI with vanishing complex phase scenario. As for Negativity, the maximum value appears in the low energy for all scenarios. For the energy range [1.52][1.5-2] GeV, the Negativity values at $\delta_{CP}$= 0 and 212, differ across all three scenarios, which are SO, NSI with vanishing phase, and NSI with non-vanishing phase.
VI Conclusions
@@
In this study, we investigated the impact of off-diagonal non-standard interaction (NSI) parameters on quantum entanglement within the three-flavor neutrino oscillation framework. By reformulating three commonly used entanglement measures; Entanglement of Formation (EOF), Concurrence, and Negativity in terms of oscillation probabilities, we analyzed how quantum correlations are modified by the presence of the off-diagonal NSI parameters $\left|\epsilon_{e\mu}\right|$, $\left|\epsilon_{e\tau}\right|$, and $\left|\epsilon_{\mu\tau}\right|$, including the influence of their complex phases. We illustrate our results for two representative choices of the $\delta_{CP}$: a CP-conserving value ($\delta_{CP}=0^{\circ}$) and the current best-fit value ($\delta_{CP}=212^{\circ}$) allowed by global analyses. However, depending on the true value of the $\delta_{CP}$, if it is determined precisely in future, one can conclude the NSI effect in different quantum measures.
@@
Focusing on the DUNE experimental setup, we studied how both the magnitude and phase of these parameters affect the energy-dependent behavior of the entanglement measures. At the oscillation maximum, pronounced deviations are observed in all quantum correlation measures for all considered NSI parameters. Among the three measures, Negativity consistently shows the most pronounced sensitivity to NSI, clearly distinguishing scenarios with and without NSI non-zero phases from standard oscillations, at higher energies. Moreover, we found that the entanglement behavior is predominantly governed by the oscillation (or survival) probabilities: the appearance channel plays the dominant role when $\epsilon_{e\mu}$ or $\epsilon_{e\tau}$ is considered individually, while the disappearance channel becomes central in the case of $\epsilon_{\mu\tau}$.
As event rates are derived from the oscillation probabilities by incorporating the source flux, interaction cross-sections, and detector effects, the overall structure and trends of the event rates closely mirror those of the underlying probabilities.
A detailed analysis of the entanglement measures in the plane of neutrino energy, $E$ and CP-violating phase, $\delta_{CP}$ highlights the distinct behavior of each measure under the influence of different off-diagonal NSI parameters. In all three NSI scenarios $\left|\epsilon_{e\mu}\right|$, $\left|\epsilon_{e\tau}\right|$, and $\left|\epsilon_{\mu\tau}\right|$ with their associated complex phases, Concurrence shows a broadly similar structure, with peaks in a specific energy ranges, and for all the value of the CP-violating phase $\delta_{CP}$. The behavior of EOF is different from that of Concurrence for all three NSI $\left|\epsilon_{e\mu}\right|$, $\left|\epsilon_{e\tau}\right|$, and $\left|\epsilon_{\mu\tau}\right|$ with their associated phase cases. The NSI $\epsilon_{e\mu}$ scenario for EOF is observed to be inconsistent across all CP violation phases. However, for the $\epsilon_{e\tau}$ and $\epsilon_{\mu\tau}$ NSI scenario, EOF displays a clear distinction between the SO and NSI with and without a complex phase $\phi_{e\tau}$, along with a noticeable dependence on $\delta_{CP}$. Negativity demonstrates a stronger sensitivity to both energy and the CP-violating phase, exhibiting sharp variations in specific regions of the $E$ - $\delta_{CP}$ parameter space for all three NSI scenarios.